badboy:

There is an increase in the amount of work completed; all other factors have remained the same. As a result, the amount of power will

Ultrilliam:

It seems the question was cut off, can you post the full question here in the comments?

Ultrilliam:

Or perhaps the options?

badboy:

decrease. decrease, then increase. increase. remain the same.

Ultrilliam:

questioncove.com/updates/4fd256bce4b057e7d2216cbd Looks like the question was asked before actually, this should help ^_^

Hero:

@badboy are you familiar with the work-power formula?

badboy:

yes

Hero:

\(\textbf{Power} = \dfrac{\text{Work}}{\text{Time}}\) right?

badboy:

yea that works thanks

Hero:

Okay, so let's just abbreviate things: \(P = \dfrac{W}{T}\)

Hero:

Suppose suddenly the amount of work doubles, In other words, we just multiplied one side of the equation by two: \(P = \dfrac{2W}{T}\). What do we need to do to P?